OpenStudy (anonymous):

task 2 a. computer capacity is often measured in bits and bytes. a bits a mall unite which is a 1 or 0, in te computer's memory. a byte is 2^3 bits. a megabyte (mb) is 2^20 bytes. how many bits in a megabyte?

ganeshie8 (ganeshie8):

1 megabyte = 2^20 bytes

ganeshie8 (ganeshie8):

1 byte = 2^3 bits

ganeshie8 (ganeshie8):

so...

ganeshie8 (ganeshie8):

1 megabyte = 2^20 bytes = 2^20 * 2^3 bits = 2^(20+3) bits = ?

OpenStudy (anonymous):

oh ok thanks

ganeshie8 (ganeshie8):

np :)